# Invariant subspaces for finite index shifts in Hardy spaces

### v2

@@ Abstract.
Let $\mathbb{H}$ be the finite direct sums of $H^{2}(\mathbb{D})$. In this paper, we give a characterization of the closed subspaces of $\mathbb{H}$ which are invariant under the shift, thus obtaining a concrete Beurling-type theorem for the finite index shift. This characterization presents any such a subspace as the finite intersection, up to an inner function, of pre-images of a closed shift-invariant subspace of $H^{2}(\mathbb{D})$ under “determinantal operators” from $\mathbb{H}$ to $H^{2}(\mathbb{D})$, that is, continuous linear operators which intertwine the shifts and appear as determinants of matrices with entries given by bounded holomorphic functions. With simple algebraic manipulations we provide a direct proof that every invariant closed subspace of codimension at least two sits into a non-trivial closed invariant subspace. As a consequence every bounded linear operator with finite defect has a nontrivial closed invariant subspace.
@@
Second author is partially supported by Plan Nacional I+D grants no. PID2019-105979GB-I00 and PID2022-137294NB-I00, Spain.
The authors also acknowledge support from “Severo Ochoa Programme for Centres of Excellence in R&D” (CEX2019-000904-S) of the Ministry of Economy and Competitiveness of Spain and by the European Regional Development Fund and from the Spanish National Research Council, through the “Ayuda extraordinaria a Centros de Excelencia Severo Ochoa” (20205CEX001).
@@ 1. Introduction
Let $H^{2}(\mathbb{D})$ be the classical Hardy space on the unit disc $\mathbb{D}$. Beurling’s theorem [5, 6] states that every closed subspace of $H^{2}(\mathbb{D})$ invariant under the shift are of the form $\phi H^{2}(\mathbb{D})$, where $\phi$ is either identically 00 or an inner function.
@@
The importance of understanding the lattice of invariant subspaces for the shift operators, besides its intrinsic interest, lies in the universality property of their adjoints (see, e.g., [18, Section 1.5]). Specifically, every continuous endomorphism of defect $\delta$ in a Hilbert space can be modeled through the restriction to an invariant subspace of any backward shift of index at least $\delta$. From Beurling’s theorem, it is straightforward to understand the lattice of invariant subspaces of the shift of index one. Via the universality property, this provides an affirmative solution to the invariant subspace problem for bounded linear operators with defect $\leq 1$.
@@
In the previous statement, it might happen that either $\mathcal{Q}_{\varphi}=\mathbb{H}$ or $\mathcal{Q}_{0}=\mathbb{H}$ (or both, in case $N=\mathbb{H}$). Note also that, in case $d=1$, the previous theorem reduces to the classical Beurling’s Theorem. Indeed, in that case, the only determinantal operator is the identity opeartor $L(f_{1}{\sf e}_{1})=f_{1}$ and hence the determinantal bricks are of the form $Q_{\varphi}=\varphi H^{2}(\mathbb{D})$ and $Q_{0}=\{0\}$.
@@
As we stated in the introduction, Theorem 1.2 allows to give an affirmative answer to the invariant subspace problem for finite defect operators (for an account history and circle of ideas related to the subspace invariant problem see, e.g., [8, 10]). Namely, let $H$ be a separable Hilbert space, and let $T:H\to H$ be a bounded linear operator. The defect of $T$ is
@@
Roughly speaking, $\delta(T)$ “measures” how much far from an isometry $T$ is. By the universality of the shifts (see, e.g., [18, Section 1.5]), every operator $T:H\to H$ with $\delta(T)\leq d$ is unitarily equivalent to the restriction of the backward shift $S^{\ast}$ to some of its invariant closed subspaces. By Theorem 1.2, the backward shift $S^{\ast}:\mathbb{H}\to\mathbb{H}$ does not have non-trivial minimal invariant closed subspaces of dimension greater than $1$. Hence, we have
@@
Let $H$ be a separable Hilbert space. Let $T:H\to H$ be a bounded linear operator such that $\delta(T)<\infty$. Then there exists a closed subspace $M\subsetneq H$, $M\neq\{0\}$ such that $T(M)\subseteq M$.
@@ 2. Preliminaries
The elements of $\mathbb{H}$ are given by $\sum_{j=1}^{d}f_{j}{\sf e}_{j}$ where $f_{j}\in H^{2}(\mathbb{D})$, $j=1,\ldots,d$. The vector space $\mathbb{H}$ is a (separable) Hilber space with Hermitian product given by
@@
As customary, if $X$ is a Hilbert space and $T:X\to X$ is a bounded linear operator, we say that a closed subspace $Y\subset X$ is $T$-invariant is $T(Y)\subseteq Y$.
@@
Let $\phi_{1},\ldots,\phi_{m}$ be inner functions. Let $\phi$ be the (inner function) greatest common divisor of $\phi_{1},\ldots,\phi_{m}$. Then there exist $m$ sequences $\{h^{j}_{n}\}_{n\in\mathbb{N}}\subset H^{2}(\mathbb{D})$, $j=1,\ldots,m$ such that $\{\sum_{j=1}^{m}h^{j}_{n}\phi_{j}\}$ converges to $\phi$ in $H^{2}(\mathbb{D})$.
@@ 3. Closed $S$-invariant spaces associated to matrices
[…]
@@ 3.2. Case $\det A\equiv 0$
Note that if $A\in\mathcal{M}_{\infty}(d\times d)$ then $A$ has rank $d$ if and only if $\det A\not\equiv 0$. On the other hand, by Kronecker’s theorem, the rank of $A$ is $k<d$ if and only if there exists a $k\times k$ matrix $B$ obtained by removing $d-k$ rows and $d-k$ columns from $A$ such that $\det B\not\equiv 0$, and all matrices containing $B$ and obtained from $A$ by removing $d-k-1$ rows and $d-k-1$ columns have determinants identically zero.
@@
Clearly, $A$ is the zero matrix (that is the $d\times d$ matrix whose entries are the identically 00 function) if and only if $\hbox{rank}(A)=0$. On the other hand, if $A$ is not the zero matrix, then $A=\phi_{A}\hat{A}$ and, since $\phi_{A}$ has at most a discrete set of zeros in $\mathbb{D}$, it is clear by the previous observation that $\hbox{rank}(\hat{A})=\hbox{rank}(A)$.
@@
The set of all indices $S$ which satisfy the previous condition with respect to $J$ is denoted by $\mathcal{J}_{J}(\hat{A})$.
@@
Let $A\in\mathcal{M}_{\infty}(d\times d)$ and assume $\det A\equiv 0$ and $\hbox{rank}(A)=k\in\{1,\ldots,d-1\}$. Let $J$ be a good multi-index of $\hat{A}$. We let $\varphi_{J,A}$ to be the (inner function) greatest common divisor of the inner factors of $\det\hat{A}^{S,J}$, when $S$ varies in $\mathcal{J}_{J}(\hat{A})$. That is,
$\varphi_{J,A}:=\hbox{g.c.d}\{I(\det\hat{A}^{S,J}):S\in\mathcal{J}_{J}(\hat{A})\}.$
@@ 4. The spaces $\mathcal{N}_{A}$ and $\mathcal{N}_{J,A}$ are not maximal for the shift
Let $k\in\{1,\ldots,d-1\}$ be the rank of $A$–and hence of $\hat{A}$. In order to simplify notation, and without loss of generality, we can assume that $J=\{1,\ldots,k\}$ is a good multi-index of $\hat{A}$ and that $J\in\mathcal{J}_{J}(\hat{A})$. Thus, $\det\hat{A}^{J,J}\not\equiv 0$.
@@
In particular, if $\theta\not\equiv 0$, by (3.3), $\theta{\sf e}_{d}\not\in\mathcal{R}(A)$ and, hence $\mathcal{R}(A)\neq\mathbb{H}$. Note that this also implies that $\dim(\mathbb{H}/\mathcal{R}(A))=\infty$—hence, $\dim(\mathcal{R}(A))^{\perp}=\infty$ and hence $\dim(\mathcal{N}_{A,J})^{\perp}=\infty$—because $[\theta{\sf e}_{d}]=[0]$ in $\mathbb{H}/\mathcal{R}(A)$ if and only if $\theta\equiv 0$ and thus $H^{2}(\mathbb{D})\ni\theta\mapsto[\theta{\sf e}_{d}]\in\mathbb{H}/\mathcal{R}(A)$ is a linear injective operator.
@@
Since $k=d-1$, we have that $J=\{1,\ldots,d-1\}$ (and we are assuming $\det\hat{A}^{J,J}\not\equiv 0$). Let $\theta$ be a non-invertible inner function coprime with the inner part $I(\det\hat{A}^{J,J})$ of $\det\hat{A}^{J,J}$ (for instance, if $I(\det\hat{A}^{J,J})(z_{0})\neq 0$ and for some $z_{0}\in\mathbb{D}$, we can take $\theta(z)=\frac{z_{0}-z}{1-\overline{z_{0}}z}$, $z\in\mathbb{D}$). Let
@@ 5. The Beurling-Lax matrix of an invariant subspace
:=$\mathbb{H}_{\infty}:=\{f_{1}{\sf e}_{1}+\ldots+f_{d}{\sf e}_{d}:f_{j}\in H^{\infty}(\mathbb{D}),j=1,\ldots,d\}.$
@@ 6. The structure theorem
Now, we claim that $W\subseteq T(A)$. To this aim, because of the (multi-)linearity of the determinant, it is enough to show that $\underline{\hat{a}}_{m}:=\hat{a}_{m1}{\sf e}_{1}+\ldots+\hat{a}_{md}{\sf e}_{d}\in T(A)$ for $m=1,\ldots,d$. So, fix $m\in\{1,\ldots,d\}$. Obviously, $\underline{\hat{a}}_{m}\in\mathcal{R}(A)$. Next, let $S=\{1\leq s_{1}<\ldots<s_{k}\leq d\}\in\mathcal{J}_{J}(\hat{A})$ and $j\in\{1,\ldots,k\}$. We need to show that $L_{\hat{A}^{S,J},j,J}(\underline{\hat{a}}_{m})\in\varphi_{J,A}H^{2}(\mathbb{D})$. Now,
@@
There are two possibilities: either the determinant is identically zero—so that $L_{\hat{A}^{S,J},j,J}(\underline{\hat{a}}_{m})\in\varphi_{J,A}H^{2}(\mathbb{D})$—or it is not identically zero. In the latter case it means that, up to reordering in increasing way, the multi-index $\{s_{1},\ldots,s_{j-1},m,s_{j+1},\ldots,s_{k}\}\in\mathcal{J}_{J}(\hat{A})$. Hence, by definition of $\varphi_{J,A}$, it follows that $\varphi_{J,A}$ divides the inner part of $L_{\hat{A}^{S,J},j,J}(\underline{\hat{a}}_{m})$, that is, $L_{\hat{A}^{S,J},j,J}(\underline{\hat{a}}_{m})\in\varphi_{J,A}H^{2}(\mathbb{D})$.
In order to show that $T(A)\subseteq W$, let $S=\{1\leq s_{1}<\ldots<s_{k}\leq d\}\in\mathcal{J}_{J}(\hat{A})$ and let
@@
On the other hand, note that $P$ is closed and $S$-invariant (since it is the closure of the sum of $S$-invariant spaces). Also, note that $\underline{\hat{a}_{m}}:=\hat{a}_{m1}{\sf e}_{1}+\ldots+\hat{a}_{md}{\sf e}_{d}$ belongs to some $T_{S}$ (indeed, if $m$ is part of some multi-index $S\in\mathcal{J}_{J}(\hat{A})$ then $\underline{\hat{a}_{m}}\in T_{S}$, while, if $m$ does not belong to any multi-index $S\in\mathcal{J}_{J}(\hat{A})$ then $L_{\hat{A}^{S,J},j,J}(\underline{\hat{a}_{m}})\equiv 0$ for all $S\in\mathcal{J}_{J}(\hat{A})$ and $j\in\{1,\ldots,k\}$, so that $\underline{\hat{a}_{m}}\in T_{S}$ for all $S\in\mathcal{J}_{J}(\hat{A})$). Thus, $W\subseteq P$ and therefore
@@
By Corollary 5.2, it is enough to show that $T(A)\cap\mathbb{H}_{\infty}\subseteq P$. To this aim, let $F:=f_{1}{\sf e}_{1}+\ldots+f_{d}{\sf e}_{d}\in T(A)\cap\mathbb{H}_{\infty}$. For every $S\in\mathcal{J}_{J}(\hat{A})$ let $\theta_{S}$ be the inner function (unique up to multiplication by a unimodular constant) such that $\varphi_{S}=\varphi_{J,A}\theta_{S}$. Note that the (inner function) greatest common divisor of $\{\theta_{S}\}_{S\in\mathcal{J}_{J}(\hat{A})}$ is $1$ (for otherwise $\varphi_{J,A}$ would not be the greatest common divisor of the $\varphi_{S}$’s).
Fix $S\in\mathcal{J}_{J}(\hat{A})$ and $j\in\{1,\ldots,k\}$. Since by hypothesis $L_{\hat{A}^{S,J},j,J}(F)\in\varphi_{J,A}H^{2}(\mathbb{D})$, it follows by the (multi-)linearity of the determinant that
@@
Thus, $\theta_{S}F\in P$ and, in particular, since $F\in\mathbb{H}_{\infty}$, we have $h\theta_{S}F\in P$ for all $h\in H^{2}(\mathbb{D})$. Hence, if for each $S\in\mathcal{J}_{J}(\hat{A})$ we take a function $h^{S}\in H^{2}(\mathbb{D})$, we have

### v3

@@ Abstract.
Let $\mathbb{H}$ be the finite direct sums of $H^{2}(\mathbb{D})$. In this paper, we give a characterization of the closed subspaces of $\mathbb{H}$ which are invariant under the shift, thus obtaining a concrete Beurling-type theorem for the finite index shift. This characterization presents any such a subspace as the finite intersection, up to an inner function, of pre-images of a closed shift-invariant subspace of $H^{2}(\mathbb{D})$ under “determinantal operators” from $\mathbb{H}$ to $H^{2}(\mathbb{D})$, that is, continuous linear operators which intertwine the shifts and appear as determinants of matrices with entries given by bounded holomorphic functions. With simple algebraic manipulations we provide a direct proof that every invariant closed subspace of codimension at least two sits into a non-trivial closed invariant subspace. As a consequence every contraction with finite defect has a nontrivial closed invariant subspace.
@@
Second author is partially supported by Plan Nacional I+D grant no. PID2022-137294NB-I00, Spain.
The authors also acknowledge support from “Severo Ochoa Programme for Centres of Excellence in R&D” (CEX2019-000904-S & CEX2023- 001347) of the Ministry of Economy and Competitiveness of Spain and by the European Regional Development Fund and from the Spanish National Research Council, through the “Ayuda extraordinaria a Centros de Excelencia Severo Ochoa” (20205CEX001).
@@ 1. Introduction
Let $H^{2}(\mathbb{D})$ be the classical Hardy space on the unit disc $\mathbb{D}$. Beurling’s theorem [5, 6] states that every closed subspace of $H^{2}(\mathbb{D})$ invariant under the shift are of the form $\phi H^{2}(\mathbb{D})$, where $\phi$ is either identically 0 or an inner function.
@@
The importance of understanding the lattice of invariant subspaces for the shift operators, besides its intrinsic interest, lies in the universality property of their adjoints (see, e.g., [18, Section 1.5]). Specifically, every continuous endomorphism of defect $\delta$ in a Hilbert space can be modeled (up to a constant) through the restriction to an invariant subspace of any backward shift of index at least $\delta$. From Beurling’s theorem, it is straightforward to understand the lattice of invariant subspaces of the shift of index one. Via the universality property, this provides an affirmative solution to the invariant subspace problem for bounded linear operators with defect $\leq 1$.
@@
In the previous statement, it might happen that either $\mathcal{Q}_{\varphi}=\mathbb{H}$ or $\mathcal{Q}_{0}=\mathbb{H}$ (or both, in case $N=\mathbb{H}$). Note also that, in case $d=1$, the previous theorem reduces to the classical Beurling’s Theorem. Indeed, in that case, the only determinantal operator is the identity operator $L(f_{1}{\sf e}_{1})=f_{1}$ and hence the determinantal bricks are of the form $Q_{\varphi}=\varphi H^{2}(\mathbb{D})$ and $Q_{0}=\{0\}$.
@@
As we stated in the introduction, Theorem 1.2 allows to give an affirmative answer to the invariant subspace problem for contractions with finite defect (for an account history and circle of ideas related to the subspace invariant problem see, e.g., [8, 10]). Namely, let $H$ be a separable Hilbert space, and let $T:H\to H$ be a contraction, namely, $\|T\|\leq 1$. The defect of $T$ is
@@
Roughly speaking, $\delta(T)$ “measures” how much far from an isometry $T$ is. By the universality of the shifts (see, e.g., [18, Section 1.5]), every contraction $T:H\to H$ with $\delta(T)\leq d$ is unitarily equivalent to the restriction of the backward shift $S^{\ast}$ to some of its invariant closed subspaces. By Theorem 1.2, the backward shift $S^{\ast}:\mathbb{H}\to\mathbb{H}$ does not have non-trivial minimal invariant closed subspaces of dimension greater than $1$. Hence, we have
@@
Let $H$ be a separable Hilbert space. Let $T:H\to H$ be a contraction such that $\delta(T)<\infty$. Then there exists a closed subspace $M\subsetneq H$, $M\neq\{0\}$ such that $T(M)\subseteq M$.
@@ Acknowledgements
Conflict of interest and Data Availability
On behalf of all authors, the corresponding author states that there is no conflict of interest.
No datasets were generated or analysed during the current study.
@@ 2. Preliminaries
The elements of $\mathbb{H}$ are given by $\sum_{j=1}^{d}f_{j}{\sf e}_{j}$ where $f_{j}\in H^{2}(\mathbb{D})$, $j=1,\ldots,d$. The vector space $\mathbb{H}$ is a (separable) Hilbert space with Hermitian product given by
@@
As customary, if $X$ is a Hilbert space and $T:X\to X$ is a bounded linear operator, we say that a closed subspace $Y\subset X$ is $T$-invariant if $T(Y)\subseteq Y$.
@@
Let $\phi_{1},\ldots,\phi_{m}$ be inner functions. Let $\phi$ be the (inner function) greatest common divisor of $\phi_{1},\ldots,\phi_{m}$. Then there exist $m$ sequences $\{h^{j}_{n}\}_{n\in\mathbb{N}}\subset H^{2}(\mathbb{D})$, $j=1,\ldots,m,$ such that $\{\sum_{j=1}^{m}h^{j}_{n}\phi_{j}\}$ converges to $\phi$ in $H^{2}(\mathbb{D})$.
@@ 3. Closed $S$-invariant spaces associated to matrices
Let $A\in\mathcal{M}_{\infty}(d\times d)$. Let $\mathcal{A}(A)=\{a_{jk}:a_{jk}\not\equiv 0\}$. If $A$ is not the zero matrix—hence $\mathcal{A}(A)\neq\emptyset$—we denote by $\phi_{A}$ the (inner function) greatest common divisor of $\{I(a_{jk}):a_{jk}\in\mathcal{A}(A)\}$. We call $\phi_{A}$ the inner greatest common divisor of $A$.
@@ 3.2. Case $\det A\equiv 0$
Note that if $A\in\mathcal{M}_{\infty}(d\times d)$ then $A$ has rank $d$ if and only if $\det A\not\equiv 0$. On the other hand, by Kronecker’s theorem, the rank of $A$ is $k<d$ if and only if there exists a $k\times k$ matrix $B$ obtained by removing dkd-k rows and dkd-k columns from $A$ such that $\det B\not\equiv 0$, and all matrices containing $B$ and obtained from $A$ by removing dk1d-k-1 rows and dk1d-k-1 columns have determinants identically zero.
@@
Clearly, $A$ is the zero matrix (that is the $d\times d$ matrix whose entries are the identically 0 function) if and only if $\hbox{rank}(A)=0$. On the other hand, if $A$ is not the zero matrix, then $A=\phi_{A}\hat{A}$ and, since $\phi_{A}$ has at most a discrete set of zeros in $\mathbb{D}$, it is clear by the previous observation that $\hbox{rank}(\hat{A})=\hbox{rank}(A)$.
@@
[…]
@@
[…]
@@ 4. The spaces $\mathcal{N}_{A}$ and $\mathcal{N}_{J,A}$ are not maximal for the shift
[…]
@@
In particular, if $\theta\not\equiv 0$, by (3.3), $\theta{\sf e}_{d}\not\in\mathcal{R}(A)$ and, hence $\mathcal{R}(A)\neq\mathbb{H}$. Note that this also implies that $\dim(\mathbb{H}/\mathcal{R}(A))=\infty$—hence, $\dim(\mathcal{R}(A))^{\perp}=\infty$ and hence $\dim(\mathcal{N}_{A,J})^{\perp}=\infty$—because $\theta{\sf e}_{d}]=[0]$ in $\mathbb{H}/\mathcal{R}(A)$ if and only if $\theta\equiv 0$ and thus $H^{2}(\mathbb{D})\ni\theta\mapsto[\theta{\sf e}_{d}]\in\mathbb{H}/\mathcal{R}(A)$ is a linear injective operator.
@@
Since k=d1k=d-1, we have that $J=\{1,\ldots,d-1\}$ (and we are assuming $\det\hat{A}^{J,J}\not\equiv 0$). Let $\theta$ be a non-invertible inner function coprime with the inner part $I(\det\hat{A}^{J,J})$ of $\det\hat{A}^{J,J}$ (for instance, if $I(\det\hat{A}^{J,J})(z_{0})\neq 0$ and for some $z_{0}\in\mathbb{D}$, we can take $\theta(z)=\frac{z_{0}-z}{1-\overline{z_{0}}z}$, $z\in\mathbb{D}$). Let
@@ 5. The Beurling-Lax matrix of an invariant subspace
$\mathbb{H}_{\infty}:=\{f_{1}{\sf e}_{1}+\ldots+f_{d}{\sf e}_{d}:f_{j}\in H^{\infty}(\mathbb{D}),j=1,\ldots,d\}.$
@@ 6. The structure theorem
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]